Mobile Company



Mobile Company manufactures computer technology devices. Selected financial data for Mobile is presented below; use the information to answer the following questions:


Current Assets As of Dec. 31, 2010 Dec. 31, 2009
Cash and short-term investments $1,267,038 $ 616,604
Accounts Receivable (net) 490,816 665,828
Inventories 338,599 487,505
Prepaid Expenses and other current assets 292,511 291,915
Total Current Assets $2,388,964 $2,061,852
Current Liabilities
Short-term borrowings $ 25,190 $ 38,108
Current portion of long-term debt 182,295 210,090
Accounts payable 296,307 334,247
Accrued liabilities 941,912 743,999
Income taxes payable 203,049 239,793
Total Current Liabilities 1,648,753 1,566,237


Selected Income Statement Data - for the year ending December 31, 2010:
Net Sales $4,885,340
Cost of Goods Sold 2,542,353
Operating Income 733,541
Net Income 230,101


Selected Statement of Cash Flow Data - for the year ending December 31, 2010:
Cash Flows from Operations $1,156,084



Refer to the information for Mobile Company. Mobile's quick ratio changed by what percentage from 2009 to 2010?

a. 25%
b. 30%
c. 82%
Answer #1

2010

Quick or acid test ratio = (current assets-inventory)/current liablities
Quick or acid test ratio = (2388964-338599)/1648573
Quick or acid test ratio = 1.24

2009

Quick or acid test ratio = (current assets-inventory)/current liablities
Quick or acid test ratio = (2061852-487505)/1566237
Quick or acid test ratio = 1.01

%age change = (ending value/beginning value-1) = 1.24/1.01-1 = 25%